OpenStudy (anonymous):

Is this correct? Giving medal Use a clock or stopwatch to take your pulse rate while sitting relaxed. Record the number of beats for thirty seconds. Now, walk out your front door and run around your house, then come back inside. Take your pulse again and record the number of beats for thirty seconds. What is the percent increase in your pulse rate? Show your work. Round your percent to the nearest whole number.

OpenStudy (anonymous):

It's raining outside, so I'm going to guess my running amount of beats. Relax=70 Running=105 I did 105-70=35 35/70=.5 .5 * 100= 50 So a 50% change in beats.

OpenStudy (anonymous):

Yup that's right, you would have an increase of 50% from running around, but that does seem a little high, did you do it for a minute instead?

OpenStudy (anonymous):

I didn't actually run outside, I don't really know much about my beats per minute for running, and I can't go outside, it's raining xD So I'm sure my teacher will understand if I just say those numbers are guessed. It still shows I know percent change doesn't it?

OpenStudy (anonymous):

Yes it shows you know percent change.